Find the common ratio
2 12 72 432

Answer by jim_thompson5910(35256)

To find the common ratio, simply divide a selected term by its previous term.

r = 12/2 = 6

r = 72/12 = 6

r = 432/72 = 6


So the common ratio is r = 6.